OpenStudy (anonymous):

3x + 2y = 4 2x - y = 5 This system of equations has no solution. has one solution. correct? is coincident.

OpenStudy (whpalmer4):

Multiply the second equation by 2, and add the two equations together. What do you get?

OpenStudy (anonymous):

no solution?

OpenStudy (whpalmer4):

what do you get after doing the steps I described?

OpenStudy (anonymous):

i dont know how to do it

OpenStudy (whpalmer4):

\[3x+2y=4\]\[2x-y=5\]Multiply the second equation by 2: \[2*2x-2*y=2*5\]\[4x-2y=10\] Add the two equations together: \[3x+4x + 2y - 2y = 4 + 10\]\[7x + 0y = 14\]Solve that for \(x\)

OpenStudy (anonymous):

so would it be coincident?

OpenStudy (whpalmer4):

Please, do as I ask…solve that equation for \(x\).

OpenStudy (anonymous):

10x + 4x + 2y

OpenStudy (whpalmer4):

No. \[7x+0y=14\]Solve that for \(x\). You'll get a number.

OpenStudy (anonymous):

7?

OpenStudy (whpalmer4):

if x = 7, then \[7(x) + 0y =14\]\[7(7) + 0 = 14\]\[49 = 14\]Is that true?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

49 = 14?!?

OpenStudy (whpalmer4):

In other developments, red is blue, black is white, up is down, left is right, did I miss any? :-)

OpenStudy (anonymous):

no. 2

OpenStudy (whpalmer4):

Much better! If \(x=2\), what does \(y=\)?

OpenStudy (anonymous):

0

OpenStudy (whpalmer4):

No, we have to go back to one of the previous equations to find \(y\): \[3x+2y=4\]\[ 2x−y=5\]\[3(2) + 2y = 4\]\[6+2y=4\]\[2y=4-6\]\[2y=-2\]\[y=-1\] or\[2(2)-y=5\]\[4-y=5\]\[-y=1\]\[y=-1\] So \((2,-1\) is the solution to the system of linear equations. Both of those equations are equations of a line, and \((2,-1)\) is the point at which they int

OpenStudy (anonymous):

okkkk has no solution. has one solution. is coincident.

OpenStudy (whpalmer4):

How many solutions did you see? The lines crossed in exactly 1 point, right? The other cases are: parallel lines: identical slope, different y-intercepts, don't ever cross, and so there are no solutions. If we go through the solution process, we end up with something like \(0=1\). That's your indication that there are no solutions. The difference in the equation at the end is simply the difference in the y-intercepts of the lines. coincident lines: identical slope, identical y-intercepts, and so there are infinitely many solutions because any points that appears on one line also appears on the other(s). If you solve the equations as we did, you end up with something like \(0=0\). That's your indication that there are infinitely many solutions.

OpenStudy (anonymous):

what? on the 1st paragraph you say theres no solution. and on your second paragraph there are infinitely many solutions.

OpenStudy (whpalmer4):

Those are two different cases, read carefully.

OpenStudy (anonymous):

Is no solution the answer

OpenStudy (whpalmer4):

No, we found a solution, didn't we? x=2, y = -1 satisfied both equations!

OpenStudy (whpalmer4):

I then made the mistake of trying to explain all 3 possible answers to you, so that you might recognize them in the future, but apparently only confused you.